OpenStudy (anonymous):

find the equation of the locus of a point that moves so that its distance from the line 12x-5y-1=0 is always 1 unit.

OpenStudy (shubhamsrg):

that'd be a line parallel to this line at a distance of 1 unit. use the formula for distance of a point and line and equate that to 1 you should get 2 equations

OpenStudy (anonymous):

i need someone to show me all the steps

OpenStudy (anonymous):

Let the coordinates of unknown point be (h,k) Distance of (h,k) from line 12x-5y-1=0 : \[\left| \frac{ 12h-5k-1 }{ \sqrt{12^2+5^2} } \right|\] = 1 gotcha ?

OpenStudy (anonymous):

thanks mate ! :)

OpenStudy (anonymous):

Welcome bro!

OpenStudy (anonymous):

i dnt get hw u know that u will get two answers ?

OpenStudy (anonymous):

There can be two such points.. ! If u take such a point, then there is another point which is mirror image of this point if you consider the line as the mirror.

OpenStudy (anonymous):

@yasmin95